# Prism-based compensation of group delay dispersion in the components of a femtosecond laser resonator and analysis of the influence of prism configuration on laser radiation parameters

### v1

@@ III GDD IN THE LASER CRYSTAL AND ITS COMPENSATION BY A PRISM PAIR
Figure 2: Geometrical parameters used to calculate GDD compensation using a two-prism system. Dashed lines indicate ray paths prior to additional insertion of the second prism by depth $A$.
@@ IV THE INFLUENCE OF PRISM CONFIGURATION ON LASER RADIATION PARAMETERS
Figure 5: Calculated possible combinations of $\Delta\lambda$ and $\lambda_{0}$ in the mode-locking regime for $l$ = 610 mm and different values of $A$: 3.07 mm (1), 3.16 mm (2), 3.38 mm (3), 3.6 mm (4). Numbers 1–4 correspond to the experimental data in Fig.6.
@@
Figure 6: Experimental radiation spectrum in the mode-locking regime for different values of $A$: 3.07 mm (1), 3.16 mm (2), 3.38 mm (3), 3.6 mm (4).
@@ V CALCULATIONS OF GDD USING GEOMETRIC MODELING
Figure 7: Geometric modeling of GDD compensation using a two-prism system with a distance between prisms $l$. Dashed lines indicate ray path prior to additional insertion of the second prism by depth $A$.

### v2

# Corresponding author ]

@@ III GDD IN THE LASER CRYSTAL AND ITS COMPENSATION BY A PRISM PAIR
Figure 2: Geometrical parameters used to calculate GDD compensation using a two-prism system. Dashed lines indicate ray paths prior to additional insertion of the second prism by depth AA.
@@ IV THE INFLUENCE OF PRISM CONFIGURATION ON LASER RADIATION PARAMETERS
Figure 5: Calculated possible combinations of $\Delta\lambda$ and $\lambda_{0}$ in the mode-locking regime for $l$ = 610 mm and different values of AA: 3.07 mm (1), 3.16 mm (2), 3.38 mm (3), 3.6 mm (4). Numbers 1–4 correspond to the experimental data in Fig.6.
@@
Figure 6: Experimental radiation spectrum in the mode-locking regime for different values of AA: 3.07 mm (1), 3.16 mm (2), 3.38 mm (3), 3.6 mm (4).
@@ V CALCULATIONS OF GDD USING GEOMETRIC MODELING
Figure 7: Geometric modeling of GDD compensation using a two-prism system with a distance between prisms $l$. Dashed lines indicate ray path prior to additional insertion of the second prism by depth AA.